# Think Before You Drive: World Model-Inspired Multimodal Grounding for Autonomous Vehicles

### v1

@@ 5.2 Comparison to State-of-the-arts (SOTA)
Model Robustness. ThinkDeeper excels in challenging scenarios, surpassing the next best baseline (UNINEXT) by 6.4% on the Corner-case set and 9.7% on the Long-text set. Similarly, on the Long-text set, it achieves 73.18 ${IoU}_{0.5}$ (+4.38 over VLTVG), highlighting its robustness in parsing extended commands, resolving linguistic ambiguity, and grounding in multi-agent scenes. Conversely, we observe that recent LLMs (MiniGPT-v2, LLaVA-NeXT, Qwen2-VL) perform poorly on these grounding tasks, with scores lagging 15-25 points behind SOTA. This is expected, as these general-purpose models lack the inductive biases for high-precision localization. Their inability to exploit 3D spatial context and depth cues to resolve ambiguity underscores the need for our depth-aware and spatial-aware world-model design tailored to real-world autonomous driving development.
@@ Appendix D Training Loss
To train our model, we calculate loss by the prediction set $\left\{s,y\right\}$ and ground-truth set $\left\{\hat{s},\hat{y}\right\}$, with $Z_{v}$ being the output of SA-WM. To fully utilize the information from each visual block, we require the model to simultaneously calculate and fit the IoU values between each visual block and the ground-truth region, which is denoted as $y$.
Stage-1: World-Model Rollout Pretraining. In the first stage, traditional loss functions such as dice loss and cross-entropy often struggle with imbalance issues in ground truth representation due to the expansive observation areas in real traffic scenarios, numerous reference objects and samples, and a low proportion of positive objects. To address this, we draw inspiration from the field of medical imaging detection of minor lesions and integrate an adapted Tversky loss $\mathcal{L}_{tve}$ [salehi2017tversky] with Focal loss $\mathcal{L}_{foc}$ [lin2017focal]. Specifically, we initialize the value of $P$ to 1 during the calculation of $S$ in the first stage to accelerate the model’s visual-text mapping learning. This strategy leverages prior scores $S$ and the downsampled ground-truth mask vector $F_{GTmask}$, interpolated via bilinear sampling, to compute the World-Model Rollout Pretraining loss. Given elements $p_{i}$ and $g_{i}$ corresponding to the $i$-th position of $S$ and $\hat{y}$, the loss ${L}_{rol}$ is formulated as follows:
${L}_{rol}$ $=\sum_{i=1}^{N}\left[\lambda_{tve}\,\mathcal{L}_{tve}(p_{i},g_{i})+\lambda_{foc}\,\mathcal{L}_{foc}(p_{i},g_{i})\right]$ (18)
$=\lambda_{tve}-\frac{\lambda_{tve}|p_{1}\cap g_{1}|}{|p_{1}\cap g_{1}|+\alpha|p_{1}\cap g_{0}|+\beta|p_{0}\cap g_{1}|}$
$\quad\quad\quad-\lambda_{foc}\,(\alpha_{t}(1-p_{t})^{\gamma}\log(p_{t}))$
where the $\lambda_{tve}$ and $\lambda_{foc}$ are both the hyperparameters. Moreover, $p_{1}$ represents the set of scores in $S$ predicted as positive examples, while $p_{0}$ denotes negative examples. Correspondingly, $|p_{1}\cap g_{1}|$ represents true positives (TP), while $|p_{1}\cap g_{0}|$ represents false negatives (FN). By adjusting the weighting parameters $\alpha$ and $\beta$, the model dynamically modulates its focus between false positives (FP) and false negatives (FN), improving class imbalance handling performance.
For focal loss $\mathcal{L}_{foc}$, the probability $p_{t}$ is defined as:
$p_{t}=\begin{cases}p,&\text{if class }t\text{ is positive}\\ 1-p,&\text{otherwise}\end{cases}$ (19)
where $p$ is the model’s estimated probability of the class being present. $\gamma$ is a focusing parameter that reduces the weight of well-classified examples, prioritizing hard, misclassified cases. $\alpha_{t}$ is a class-specific weighting factor to further adjust the impact of positive and negative samples. This loss enhances model flexibility, ensuring robust adaptation to imbalanced classes while improving visual-text alignment.
Stage-2: Grounding Decision Supervision. In the second stage, to prevent the model from forgetting previously acquired patch-level task knowledge, we employ a multi-task learning strategy to optimize both Binary Cross-Entropy (BCE) loss $\mathcal{L}_{bce}$ and L1 loss $\mathcal{L}_{L1}$. We calculate loss ${L}_{gro}$ by the prediction set $\left\{s,y\right\}$ and ground-truth set $\left\{\hat{s},\hat{y}\right\}$, with the future latent states $S$ is the SA-WM’s output that represents the mask annotations of referred object. Formally,
[…]
@@
where $\lambda_{bce}$ and $\lambda_{L1}$ are the hyperparameters. Overall, this diversity loss term incentivizes the model to further capture cross-modal interactions and progressively produce target predictions consistent with the commander’s intent.

### v2

@@ 5.2 Comparison to State-of-the-arts (SOTA)
Model Robustness. ThinkDeeper excels in challenging scenarios, surpassing the next best baseline (UNINEXT) by 7.2% on the Corner-case set and 11.8% on the Long-text set. Specifically, on the Long-text set, it achieves 74.08 ${IoU}_{0.5}$ (+5.28 over VLTVG), highlighting its robustness in parsing extended commands, resolving linguistic ambiguity, and grounding in multi-agent scenes. Conversely, we observe that recent LLMs (MiniGPT-v2, LLaVA-NeXT, Qwen2-VL) perform poorly on these grounding tasks, with scores lagging 15-25 points behind SOTA. This is expected, as these general-purpose models lack the inductive biases for high-precision localization. Their inability to exploit 3D spatial context and depth cues to resolve ambiguity underscores the need for our depth-aware and spatial-aware world-model design tailored to real-world autonomous driving development.
@@ Appendix D Training Loss
To train our model, we define supervision over the prediction tuple $\{s,y\}$ and the ground-truth tuple $\{\hat{s},\hat{y}\}$, where $Z_{v}$ denotes the visual latent states produced by SA-WM. For each visual block, the model is required to regress the IoU between the predicted region and the annotated ground-truth region, which is denoted as $y$. This design ensures that every visual block contributes explicit grounding signals rather than relying solely on a final aggregated score.
Stage-1: World-Model Rollout Pretraining. In the first stage, the supervision is dominated by dense, patch-level observations in complex traffic scenes. Under such settings, standard region-based losses like Dice loss or vanilla cross-entropy frequently suffer from severe foreground–background imbalance, caused by large fields of view, multiple reference objects, and a relatively small proportion of positive pixels. To mitigate this, we draw inspiration from lesion detection in medical imaging and combine a Tversky loss $\mathcal{L}_{\mathrm{tve}}$ [salehi2017tversky] with Focal loss $\mathcal{L}_{\mathrm{foc}}$ [lin2017focal]. During World-Model rollouts, we initialize the prior score $S$ with uniform probability to accelerate the convergence of visual-text mapping. By leveraging bilinear interpolation on the downsampled ground-truth mask $\hat{{S}}$, we enforce a dense alignment constraint.
Unlike standard formulations that simply sum losses, we construct a weighted integration of a re-parameterized Tversky objective and a hardness-weighted Focal penalty. Let $p_{ij}\in{S}$ and $g_{ij}\in\hat{{S}}$ denote the predicted probability and binary ground truth at spatial location $(i,j)$. We define the intersection and set differences for the $k$-th class as follows:
$\mathcal{I}_{k}$ $=\sum\nolimits_{i,j}p_{ij}g_{ij},$
$\mathcal{F}_{k}^{\alpha}$ $=\alpha\sum\nolimits_{i,j}(1-g_{ij})p_{ij},\;\;\mathcal{F}_{k}^{\beta}=\beta\sum\nolimits_{i,j}(1-p_{ij})g_{ij}$ (18)
where $\epsilon$ is a smoothing factor, and $\alpha,\beta$ control the penalty magnitude for False Positives (FP) and False Negatives (FN), respectively, allowing the model to dynamically shift focus during training. Then, the rollout loss ${L}_{rol}$ is formulated as:
$\begin{split}{L}_{\text{roll}}=\lambda_{tve}\sum_{k=1}^{K}\left[1-\frac{\mathcal{I}_{k}+\epsilon}{\mathcal{I}_{k}+\mathcal{F}_{k}^{\alpha}+\mathcal{F}_{k}^{\beta}+\epsilon}\right]-\lambda_{\text{foc}}\frac{1}{N}\sum_{n=1}^{N}\mathcal{H}(p_{n},g_{n})\end{split}$ (19)
Here, the $\lambda_{tve}$ and $\lambda_{foc}$ are both the hyperparameters. Moreover, $\mathcal{H}(\cdot)$ represents the focal modulation defined as $\mathcal{H}(p_{n},g_{n})=\alpha_{t}(1-p_{t,n})^{\gamma}\log(p_{t,n})$, where $p_{t,n}$ reflects the model’s confidence in the true class. This formulation ensures that gradients are dominated by hard-to-classify examples like small distant vehicles, rather than the dominant background, effectively regularizing the rollout trajectory.
Stage-2: Grounding Decision Supervision. Upon establishing robust feature representations, the second stage focuses on precise localization. To prevent catastrophic forgetting of the patch-level priors learned in the first stage, we employ a multi-task learning strategy. We define the grounding loss ${L}_{\text{ground}}$ as a hybrid constraint optimization that simultaneously minimizes the distributional divergence and the geometric regression error. Let $\hat{\mathbf{y}}$ be the ground truth bounding box and ${S}_{lat}$ be the latent state output representing the object mask. The objective function aggregates the Binary Cross-Entropy (BCE) and L1 regression error, normalized over the batch. The supervision loss ${L}_{\text{gro}}$ can be defined as:
${L}_{\text{gro}}=\mathbb{E}_{({y},{S})\sim\mathcal{D}}\left[\lambda_{\text{cls}}\cdot\Psi_{\text{BCE}}({y},\hat{{y}})+\lambda_{\text{reg}}\cdot\|{S}_{lat}-\hat{{S}}\|_{1}\right]$ (20)
@@
where $\Psi_{\text{BCE}}$ is the binary cross-entropy operator, and $\|\cdot\|_{1}$ imposes a sparsity-inducing L1 penalty on the mask prediction. The hyperparameters $\lambda_{\text{cls}}$ and $\lambda_{\text{reg}}$ balance the trade-off between semantic classification accuracy and geometric precision. Overall, this diversity loss term incentivizes the model to capture cross-modal interactions and produce target predictions consistent with the commander’s intent.